Find the 2 real numbers whose difference is 12 and whose product is as small as possible. Justify your answer using the 1st or 2nd derivative test.
Please explain
Thank you

Answer by fractalier(6550)

We need to minimize
f(x)=x(12+x)=12x + x^2
Set the derivative equal to zero and get
f'(x) = 12 + 2x = 0
and x = -6
The second derivative is
f''(x) = 2
so you know it is a minimum.
Thus the two numbers are
-6 and 6, and the product is -36.